*Factor out the greatest common factor.
16y^5 - 24y^4 -40y^3
5a(a + 3b) + 4(a+3b)
*Factor
X^3 +8X^2 + 2x+16
2bc + 4b -5c-10
*Factor each polynomial
X^2 + 12X + 11
*Factor out the greatest common factor from the terms of the trinomial. Then factor the remaining trinomial.
2x + 30x +52

Answer by stanbon(75887)

Factor out the greatest common factor.
16y^5 - 24y^4 -40y^3
= 8y^3(2y^2-3y-5)
= 8y^3(2y^2-5y+2y-5)
= 8y^3(y(2y-5)+(2y-5)
= 8y^3(2y-5)(y+1)
--------------------------------
5a(a + 3b) + 4(a+3b)
= (a+3b)(5a+4)
--------------------------------
*Factor
X^3 +8X^2 + 2x+16
x^2(x+8)+2(x+8
= (x+8)(x^2+2)
------------------------------------
2bc + 4b -5c-10
= 2b(c+2)-5(c+2)
= (c+2)(2b-5)
==========================
*Factor each polynomial
X^2 + 12X + 11
= (x+11)(x+1)
-------------------------
*Factor out the greatest common factor from the terms of the trinomial. Then factor the remaining trinomial.
2x^2 + 30x +52
= 2(x^2+15x+26)
= 2(x^2+13x+2x+26)
= 2(x(x+13)+2(x+13))
= 2(x+13)(x+2)
====================
Cheers,
Stan H.
====================